# Adaptation of Agentic AI

### v2

@@ 8.2 Continual Adaptation
While our discussion has so far centered on agent adaptation mechanisms such as A1 and A2, these methods still assume a fixed task distribution and are typically instantiated on a single downstream task at a time. In contrast, real-world deployments involve non-stationary task distributions, where tasks, tools, and user needs evolve over time, making isolated, one-off adaptations prone to Catastrophic Forgetting (CF). This calls for Self-Evolving Agents that continuously update their behaviors, tools, and memories in open and dynamic environments. Continual Learning (CL) [235, 236, 237, 238] provides a natural foundation for this goal, as it studies how models learn from non-stationary task streams while retaining prior knowledge. We therefore revisit CL techniques that can serve as concrete mechanisms for Self-Evolving Agents and organize them into the following two categories.
@@
Figure 10: An illustrative example of continual adaptation.
Notably, this challenge of continual adaptation becomes especially pronounced in domains with strong execution-based supervision signals, such as those enabled by reinforcement learning with verifiable rewards (RLVR). As discussed in §4.1.2, environments like formal theorem proving provide reliable, tool-execution–signaled feedback for learning multi-step behaviors. At the same time, these domains often evolve structurally over time—for example, through expanding formal math libraries and large, actively maintained formalization projects—making them representative testbeds for continual agent adaptation. Instead of repeatedly retraining the core agent, many prover agent systems adapt to expanding libraries by updating premise retrieval indices, tactic databases, or proof-state memories, allowing agents to exploit newly introduced lemmas without rewriting the entire policy [73, 259]. Such low-resource adaptation complements RLVR-style training by isolating long-term knowledge growth from short-term policy optimization, and again aligns with our notion of T2 adaptation.
Taken together, these two lines of work highlight complementary trade-offs for building Self-Evolving Agents. Within the dynamic A1/A2 paradigm, recent results [260] show that not all parameter-update schemes forget equally: RL with a reverse-KL objective and on-policy data can achieve comparable or better performance than SFT while exhibiting substantially less forgetting, suggesting that on-policy data streams can act as an intrinsic CL mechanism for continual agent adaptation. Yet such methods still rewrite a shared set of parameters, so forgetting and interference are mitigated rather than structurally removed. The evolving T2 paradigm tackles CF at the architectural level by freezing the core agent and encapsulating new capabilities in external, independently trained tools or subagents, which avoids the clobbering of a monolithic parameter space. Looking ahead, a promising direction is to systematically integrate these two perspectives, using CL-aware parameter updates where they are most effective while shifting as much long-term adaptation as possible into T2-style modular tools and external memories.
@@ 8.3 Safe Adaptation
The transition from static foundation models to adaptive agentic systems marks a fundamental inflection point in AI safety. While traditional safety paradigms focus on the alignment of frozen weights, adaptation mechanisms, specifically on-policy optimization (A1) and outcome-driven tool tuning (T2), introduce dynamic threat vectors characterized by autonomous risk-taking and adversarial co-evolution [261]. We categorize these emerging risks into two primary failure modes: Unsafe Exploration, arising from stochastic trial-and-error, and Parasitic Adaptation, arising from exploitative optimization loops.
@@
  • The Reward-Safety Gap: In frameworks like RLEF [262] or DeepRetrieval [21], rewards are typically sparse and binary (e.g., task completion). This creates a feedback vacuum for intermediate actions, encouraging agents to maximize efficacy regardless of collateral damage (e.g., deleting system files to free space) [265].
@@
  • Type B: Adversarial Tooling (The Tool as Parasite): In T2 ecosystems utilizing protocols like MCP [129], tools can evolve to exploit the agent. A compromised or parasitic tool may return prompt-injected data that hijacks the agent’s reasoning (the “Confused Deputy” problem), forcing the agent to exfiltrate sensitive data under the guise of standard tool use [273, 274].
@@
Addressing these risks requires moving beyond scalar rewards toward robust specification. A straightforward yet effective mitigation for Security Risk 1 is to introduce a safety-check layer before the agent’s input reaches the tool (Figure 11). This gate ensures that any anomalous or unsafe behaviors are intercepted and filtered out prior to execution. More sophisticated solutions include: Constrained Policy Optimization [276, 277, 278] and safety shields project agent actions onto verified safe sets to prevent catastrophic exploration. Verifiable Rewards [279, 58] replace opaque preference models with programmatic outcome verification (e.g., unit tests, proofs) to reduce sycophancy. Specification Self-Correction [280] allows agents to dynamically critique and refine reward functions at inference time to detect gaming. Finally, Proof-of-Use [281] frameworks enforce causal links between retrieved evidence and generated answers, preventing tool-use hallucination.
@@ 8.4 Efficient Adaptation
Efficient adaptation in agentic AI presents several important opportunities. Current systems typically rely on large-scale GPU clusters for fine-tuning or policy refinement, which limits accessibility and personalization. Shifting adaptation toward resource-constrained settings could enable agents to learn efficiently on mobile devices, edge hardware, or other low-power environments. This would not only broaden the applicability of agentic AI but also allow continual adaptation directly on user devices while preserving privacy. Moreover, learning close to the interactions that generate the training signal could reduce latency between experience and model update, effectively blurring the distinction between inference and training. Within this context, we identify several concrete opportunities:
Parameter-Efficient Adaptation: Techniques such as Low-Rank Adaptation (LoRA) [19] and its extensions [282, 283, 284, 285, 286, 287, 288] allow large models to adapt to new tasks by updating only a small subset of weights, significantly reducing memory and computational requirements. Recent work LoRA Without Regrets [289] empirically demonstrates that LoRA can be effectively applied in reinforcement learning settings. They provide evidence that LoRA performs equivalently to full fine-tuning even at small ranks in RL tasks, indicating that RL often requires very low parameter capacity. This observation suggests that models can be fine-tuned on resource-constrained devices while maintaining strong RL performance. Figure 12 shows an illustrative example of this.
Quantized Adaptation: FlashRL [290] introduces a framework for accelerating reinforcement learning by performing rollout generation in lower numerical precision, such as INT8 or FP8, while preserving downstream performance. The key innovation lies in mitigating the rollout-training mismatch caused by quantization through truncated importance sampling (TIS), which stabilizes gradient estimation when the agent’s policy generating rollouts is quantized, but the training engine remains in higher precision. Empirical results demonstrate that FlashRL can achieve significant speedups in RL training without sacrificing final task performance. This work provides compelling empirical evidence that quantization, a technique widely employed in supervised model inference, can be effectively extended to agentic reinforcement learning. By enabling rollout generation at reduced numerical precision without degrading downstream performance, FlashRL demonstrates the potential for RL training across large-scale and resource-constrained environments.
@@
On-Device & Personalized Adaptation: On-device adaptation [291, 292, 293, 294, 295, 296, 297, 118] focuses on enabling agents to learn and update directly on user devices under tight computational and memory constraints. This capability has become increasingly important as modern agents operate across heterogeneous hardware, operating systems, interface designs, and interaction contexts, all of which introduce substantial variation in user behavior, application semantics, and device-specific execution patterns. A key component of on-device adaptation is personalization, which allows agents to reflect individual preferences [298, 299], maintain persistent memory [300], and adjust behavior over time [301, 302]. Recent progress in GUI agents further stresses the importance of this direction: modern GUI agents [303, 304, 305, 306, 307, 305, 308, 309] rely on strong user-specific multimodal reasoning. A promising strategy for agentic personalization is tool adaptation, where each device maintains a lightweight tool module aligned with user-specific habits and interaction patterns. Instead of modifying the full model or even parameter-efficient adapters, adapting a small tool module focuses directly on modeling personal preferences, recording relevant user history, and shaping dynamic behavioral adjustments. Since the tool module is fully decoupled from the base model, it can update locally without compromising global capabilities. Frequent and incremental updates of tool modules strengthen preference alignment, reinforce long-term memory, and support continual behavioral adaptation.
Overall, these opportunities suggest a pathway toward agentic AI that is more efficient and scalable. By combining parameter-efficient fine-tuning, quantization, and on-device adaptation, future agents can evolve continuously in close alignment with user needs and environmental constraints.
@@ 9 Conclusion
The transition from static foundation models to autonomous agentic systems marks a fundamental shift in artificial intelligence, moving from passive response generation to active and multi-step problem solving. As these systems are deployed in increasingly complex and open-ended environments, the ability to adapt to refine behavior, master new tools, and align with specific tasks has become the primary driver of reliability and performance. In this paper, we have provided a comprehensive roadmap of this landscape, introducing a unified taxonomy that organizes adaptation strategies into four distinct paradigms based on the locus of optimization and the source of the supervision signal.
@@
Our framework reveals that the design space of agentic adaptation is defined by the tension between monolithic and modular evolution. The agent centric paradigms, A1 (Tool Execution Signaled) and A2 (Agent Output Signaled), offer high parametric flexibility, allowing models to internalize tool mechanics and complex reasoning strategies through direct environmental feedback or holistic outcome evaluation. However, these approaches often incur high computational costs and risk catastrophic forgetting. Conversely, the tool centric paradigms, T1 (Agent Agnostic) and T2 (Agent Supervised), shift the burden of adaptation to the peripheral ecosystem. By treating tools and even other graduated agents as modular and optimizing components, these paradigms enable system level flexibility and significant data efficiency.
A critical insight emerging from our analysis is the symbiotic inversion represented by the T2 paradigm. Rather than treating the foundation model as the object of optimization, T2 reframes it as a stable source of supervision, training lightweight subagents such as searchers, planners, and memory curators to serve the frozen core. This architectural shift not only decouples skill acquisition from general reasoning but also paves the way for federated agentic systems that can evolve continuously without destabilizing the backbone model.
Looking forward, the advancement of agentic AI depends on the strategic integration of these paradigms rather than their isolation. Future systems will likely leverage a hybrid architecture, combining the reasoning depth of agent centric adaptation with the modular efficiency of tool centric adaptation to achieve robustness and scalability. Realizing this potential requires addressing the fundamental challenges of continual adaptation to maintain performance in dynamic streams, safe adaptation to mitigate risks such as reward hacking, and efficient adaptation to enable deployment in resource constrained environments. Ultimately, the next generation of intelligent systems will be defined not by a single monolithic model, but by the principled orchestration of stable reasoning cores supported by specialized and adaptive tools.

### v3

@@ 9.2 Continual Adaptation
While our discussion has so far centered on agent adaptation mechanisms such as A1 and A2, these methods still assume a fixed task distribution and are typically instantiated on a single downstream task at a time. In contrast, real-world deployments involve non-stationary task distributions, where tasks, tools, and user needs evolve over time, making isolated, one-off adaptations prone to Catastrophic Forgetting (CF). Self-Evolving Agents that continuously update their behaviors, tools, and memories in open and dynamic environments are therefore needed. Continual Learning (CL) [338, 339, 340, 341] provides a natural foundation for this goal, as it studies how models learn from non-stationary task streams while retaining prior knowledge. We revisit CL techniques that can serve as concrete mechanisms for Self-Evolving Agents and organize them into two categories.
@@
The challenge of continual adaptation becomes especially pronounced in domains with strong execution-based supervision signals, such as those enabled by reinforcement learning with verifiable rewards (RLVR). As discussed in §4.1.2, environments like formal theorem proving provide reliable, tool-execution-signaled feedback for learning multi-step behaviors. At the same time, these domains often evolve structurally over time (for example, through expanding formal math libraries and actively maintained formalization projects), making them representative testbeds for continual agent adaptation. Instead of repeatedly retraining the core agent, many prover agent systems adapt to expanding libraries by updating premise retrieval indices, tactic databases, or proof-state memories, allowing agents to exploit newly introduced lemmas without rewriting the entire policy [74, 362]. Such low-resource adaptation complements RLVR-style training by isolating long-term knowledge growth from short-term policy optimization, and again aligns with our notion of T2 adaptation.
These two lines of work address complementary trade-offs for building Self-Evolving Agents. Within the dynamic A1/A2 paradigm, recent results [292] show that not all parameter-update schemes forget equally: RL with a reverse-KL objective and on-policy data can achieve comparable or better performance than SFT while exhibiting substantially less forgetting, suggesting that on-policy data streams can act as an intrinsic CL mechanism for continual agent adaptation. Yet such methods still rewrite a shared set of parameters, so forgetting and interference are mitigated rather than structurally removed. The evolving T2 paradigm tackles CF at the architectural level by freezing the core agent and encapsulating new capabilities in external, independently trained tools or subagents, which avoids interference within a shared parameter space. A promising direction is to integrate these two perspectives, using CL-aware parameter updates where they are most effective while shifting as much long-term adaptation as possible into T2-style modular tools and external memories.
@@ 9.3 Safe Adaptation
Figure 11: An illustrative example of safe adaptation.
The transition from static foundation models to adaptive agentic systems introduces new challenges for AI safety. While traditional safety paradigms focus on the alignment of frozen weights, adaptation mechanisms, specifically on-policy optimization (A1) and outcome-driven tool tuning (T2), introduce dynamic threat vectors characterized by autonomous risk-taking and adversarial co-evolution [363]. We categorize these emerging risks into two primary failure modes: Unsafe Exploration, arising from stochastic trial-and-error, and Parasitic Adaptation, arising from exploitative optimization loops.
@@
  • The Reward-Safety Gap: In frameworks like RLEF [364] or DeepRetrieval [22], rewards are typically sparse and binary (e.g., task completion). The result is a feedback vacuum for intermediate actions, encouraging agents to maximize efficacy regardless of collateral damage (e.g., deleting system files to free space) [366].
@@
  • Type B: Adversarial Tooling (The Tool as Parasite): In T2 ecosystems utilizing protocols like MCP [130], tools can evolve to exploit the agent. A compromised or parasitic tool may return prompt-injected data that hijacks the agent’s reasoning (the “Confused Deputy” problem), forcing the agent to exfiltrate sensitive data under the guise of standard tool use [373, 374]. Recent empirical studies show that this risk is already material at the skill layer: Liu et al. [255] report vulnerabilities in 26.1% of marketplace skills, and Skill-Inject [254] finds up to 80% attack success rates from prompt injections embedded in skill files.
@@
Addressing these risks requires moving beyond scalar rewards toward robust specification. A direct mitigation for Security Risk 1 is a safety-check layer before the agent’s input reaches the tool (Figure 11), which intercepts and filters anomalous or unsafe behaviors prior to execution. More targeted solutions include: Constrained Policy Optimization [376, 377, 378] and safety shields project agent actions onto verified safe sets to prevent catastrophic exploration. Verifiable Rewards [379, 59] replace opaque preference models with programmatic outcome verification (e.g., unit tests, proofs) to reduce sycophancy. Specification Self-Correction [380] allows agents to dynamically critique and refine reward functions at inference time to detect gaming. Finally, Proof-of-Use [381] frameworks enforce causal links between retrieved evidence and generated answers, preventing tool-use hallucination.
@@ 9.4 Efficient Adaptation
Current agentic adaptation assumes large-scale GPU clusters for fine-tuning or policy refinement, which limits deployment to cloud settings. Efficiency matters differently in agentic contexts than in standard LLM serving: each adaptation step may involve multiple tool calls, environment interactions, and thousands of tokens, so even small per-step savings compound rapidly. We organize emerging directions by which bottleneck they address:
@@
Parameter-Efficient Adaptation: Techniques such as Low-Rank Adaptation (LoRA) [20] and its extensions [382, 383, 384, 385, 386, 387, 388] allow large models to adapt to new tasks by updating only a small subset of weights, reducing memory and computational requirements. Recent work on LoRA Without Regrets [389] shows that LoRA can be applied in reinforcement learning settings. They provide evidence that LoRA performs equivalently to full fine-tuning even at small ranks in RL tasks, indicating that RL often requires very low parameter capacity. Models can therefore be fine-tuned on resource-constrained devices while maintaining strong RL performance (Figure 12).
Quantized Adaptation: FlashRL [390] accelerates reinforcement learning by performing rollout generation in lower numerical precision (INT8 or FP8) while preserving downstream performance. Its core contribution, truncated importance sampling (TIS), stabilizes gradient estimation when the rollout policy is quantized but the training engine remains in higher precision. FlashRL achieves substantial speedups without sacrificing final task performance, providing evidence that quantization extends from supervised inference to agentic reinforcement learning.
On-Device & Personalized Adaptation: On-device adaptation [391, 392, 393, 394, 395, 396, 397, 122] enables agents to learn directly on user devices under tight computational and memory constraints. Modern agents operate across heterogeneous hardware and interaction contexts, introducing substantial variation in user behavior and device-specific execution patterns. Personalization, an important component, allows agents to reflect individual preferences [398, 399], maintain persistent memory [400], and adjust behavior over time [401, 402]. Recent GUI agents [403, 404, 405, 406, 407, 408, 409] that rely on user-specific multimodal reasoning further underscore the need for on-device personalization. A natural strategy is lightweight tool adaptation: each device maintains a small tool module aligned with user-specific habits. Because the tool module is fully decoupled from the base model, it can update locally without compromising global capabilities, supporting continual behavioral adaptation through frequent incremental updates.
These three directions target orthogonal bottlenecks—parameter count, numerical precision, and deployment location—and can be composed: a LoRA adapter quantized to INT8 and updated on-device represents the intersection of all three.
@@ 10 Conclusion
This survey organized agentic AI adaptation around a single question: when the system underperforms, should we change the agent, the tool, or both—and what signal should drive that change? The resulting four-paradigm framework (A1/A2/T1/T2) structures a rapidly growing literature along two axes—locus of optimization (agent vs. tool) and signal source (execution-grounded vs. output-evaluated)—and grounds the abstract taxonomy in three concrete mechanisms: post-training, memory, and skills.
@@
The design space is defined by the tension between monolithic and modular evolution. Agent-centric paradigms (A1, A2) offer high parametric flexibility, allowing models to internalize tool mechanics and complex reasoning strategies through direct environmental feedback (A1) or holistic outcome evaluation (A2). However, these approaches incur high computational costs and, for A2 in particular, risk degrading previously learned capabilities when the agent is adapted to new domains. The degree of forgetting is paradigm- and method-dependent: on-policy RL with reverse-KL regularization exhibits less forgetting than SFT in some settings [292], and T2-style modular adaptation avoids the problem structurally by keeping the core agent frozen.
Tool-centric paradigms (T1, T2) shift the burden of adaptation to the peripheral ecosystem. T1 provides plug-and-play reusability; T2 enables the frozen agent to supervise lightweight subagents (searchers, planners, memory curators), achieving competitive accuracy with substantially fewer training examples in case studies such as retrieval-augmented QA. We emphasize that these cross-paradigm comparisons are not yet controlled experiments: the systems differ in optimization target, backbone composition, and system architecture, so the efficiency gap cannot be attributed solely to paradigm choice. Controlled cross-paradigm benchmarks remain a critical open problem.
Several additional findings merit highlighting:
  • Signal density shapes paradigm effectiveness. A1 methods excel in domains with dense, verifiable execution feedback (code, theorem proving, SQL); A2 is necessary when only sparse outcome signals are available. The pattern holds across multiple case studies, though the evidence base for a universal claim is incomplete.
  • Evaluation is paradigm-dependent. The same system looks different under A1 metrics (tool-execution quality) versus A2 metrics (end-to-end task success), and no single metric suffices. Multi-dimensional evaluation that jointly reports accuracy, cost, safety, and adaptation dynamics is essential.
  • The graduation lifecycle bridges paradigms. Agents trained under A1/A2 can be frozen and redeployed as T1 tools, creating a cycle where agent adaptation enriches the tool ecosystem.
  • Memory and skills span the full taxonomy. Episodic buffers, reflective databases, and knowledge graphs can all be optimized as T2 tools, giving agents persistent experience without parameter updates. Skills similarly range from A1/A2-internalized tool-use patterns to external skill libraries that accumulate as reusable T1/T2 resources.
Future progress in agentic AI depends on integrating these paradigms rather than treating them in isolation. Achieving this requires advances in co-adaptation (jointly optimizing agents and tools under non-stationary dynamics), continual adaptation (maintaining performance as task distributions shift), safe adaptation (mitigating risks such as reward hacking and unsafe exploration during on-policy learning), and efficient adaptation (enabling deployment on resource-constrained devices). Progress will depend less on any single monolithic model than on principled orchestration of stable reasoning cores alongside an evolving ecosystem of specialized, adaptive tools and memory systems.